OpenStudy (anonymous):

Find the exact value of cos(arcsin(1/4)). For full credit, explain your reasoning.

OpenStudy (anonymous):

Let \(u=\sin^{-1}\left(\dfrac{1}{4}\right)\), so you have \(\sin u=\dfrac{1}{4}\).  How would you go about finding \(\cos u\)?

OpenStudy (anonymous):

A/h

OpenStudy (anonymous):

Right, which means \[\cos u=\frac{t}{4}\] What's the value of \(t\) ?

OpenStudy (anonymous):

Since you have a right triangle with two of the known sides, how would you solve for the unknown side?

OpenStudy (anonymous):

Pythagreom theorem

OpenStudy (anonymous):

Right. What do you get for \(t\) ?

OpenStudy (anonymous):

sqrt of 15

OpenStudy (anonymous):

Right again. So \(\cos u=\cos\left[\sin^{-1}\left(\dfrac{1}{4}\right)\right]=\dfrac{\sqrt{15}}{4}\)